OpenStudy (anonymous):

Suppose that a complex number u is obtained as a result of a finite number of rational operations applied to the complex numbers z1,z2,z3,..zn. Prove that the same operations applied to the complex conjugates of z1,z2,z3..zn leads to the number ü, conjugate of u.

OpenStudy (experimentx):

rational operations??

OpenStudy (anonymous):

summ , multiplication division..

OpenStudy (experimentx):

with complex number or real number?

OpenStudy (anonymous):

complex

OpenStudy (anonymous):

www.dms.uaf.edu/~rybkin/MathPhysicsLectureNotes.pdf follow this

OpenStudy (anonymous):

maybe prooving by induction for each operation separatly...

OpenStudy (experimentx):

i was thinking the same

OpenStudy (experimentx):

let a+ib be a complex number, so a-ib is it's conjugate, try adding c+id to both of them, (a+c)+(b+d)j for our number (a+c)+(d-b)j or conjugate ... Oo whats this??

OpenStudy (anonymous):

@Rohangrr in the link you posted says just check that that's true, but its not prooved

OpenStudy (anonymous):

@myko give up a medal then

OpenStudy (experimentx):

let's move on to multiplication: (a+ib) x (c+id) = (ac-bd)+i(ad+bd)) for conjugate (a+ib) x (c-id) = (ac+bd)+i(ad-bd)) ???

OpenStudy (anonymous):

I think it might be easyer by induction: Summ: supose true for n=2... Conj(z1+z2) = Conj(z1)+Conj(z2) for n=3 ...Conj(z1)Conj(z2)Conj(z3)=Conj(z1+z2) + Conj(z3) = Conj(z1+z2+z3) Similarly for multiplication

OpenStudy (anonymous):

@experimentX ??

OpenStudy (experimentx):

yeah ... let's try it!!!

OpenStudy (experimentx):

but damn ... what did i do above???

OpenStudy (anonymous):

idk

OpenStudy (experimentx):

let's prove the basic first .. (a+ib) x (c+id) = (ac-bd)+i(ad+bd)) for conjugate (a-ib) x (c+id) = (ac+bd)+i(ad-bd)) ??? let's see for conjugate (a-ib) x (c-id) = (ac-bd)-i(ad+bd)) ??? maybe you meant ... operation on conjugate bu conjugate??????

OpenStudy (experimentx):

same for (a+c)+(b+d)j for our number (a+c)-(d+b)j or conjugate ... for addition. @myko i think you operation on conjugate by conjugate

OpenStudy (anonymous):

I have to prove Conj(z1,z2,z3...zn) = Conj(z1)Conj(z2)...Conj(zn) and its true for every operation separatly I am only stragling with division by induction....

OpenStudy (anonymous):

ok, got it too

OpenStudy (experimentx):

damn ... man!!! it's what i said as before ...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i don't need a simple check, i need proof

OpenStudy (experimentx):

conj(a+b) = conj(a) + conj(b) --- we have to operate on conjugate with conjugate

OpenStudy (anonymous):

induction works fine

OpenStudy (anonymous):

thx for help

OpenStudy (experimentx):

prove this for n=2 prove this for n=3 assume this for n=n prove this for n=n+1

OpenStudy (anonymous):

thats what i did